# An Information Theoretic Approach to Machine Unlearning

### v3

@@ 1 Introduction
Regulations such as the General Data Protection Regulation (GDPR) enshrine an individual’s data autonomy rights, including the right to be forgotten. While deleting an entry from a database is relatively straightforward, removing the influence of that data from a trained model is a challenging open problem. The process of removal, referred to as unlearning, is difficult for several reasons. It is known that neural networks memorise instance-level information (Arpit et al., 2017; Zhang et al., 2021; Feldman, 2020), and it is effectively intractable to ascribe parameter changes to a specific training sample post-hoc (Kurmanji et al., 2023). At its core, unlearning is a multi-objective optimization problem with three key desiderata. An effective unlearning algorithm must remove the influence of the selected subset of data, maintain model performance on retained data, and minimise computational cost. These goals are antagonistic since inducing forgetting inevitably disrupts the model’s performance, thus balancing these objectives is key. Naïvely, one may achieve perfect forgetting by retraining a model on the training data sans the forget samples every time there is a forget request. However, this is prohibitively expensive thus violating the third desideratum.
@@
Many of the unlearning methods proposed are effective, however they make strong assumptions about the problem setting that simplifies the task considerably. Primarily, existing methods typically assume access to all, or a subset of, the training data. This data is used in different ways, such as to fine-tune the model post-forgetting (Chundawat et al., 2023a; Graves et al., 2021), or to conduct parameter importance calculations after the initial training period (Foster et al., 2023). In reality, there are many reasons why this data could be unavailable, such as cost of storage, limited duration access to datasets, or an oversight in considering machine unlearning during model development. As such, Chundawat et al. (2023b) introduces a novel problem setting for unlearning, termed zero-shot (ZS) unlearning, whereby only the data to be forgotten and the trained model are available. This is extremely challenging, since the remaining data is not available to protect model performance, and thus more delicate methods are required. Insightful treatment of the ZS scenario can be found in Chen et al. (2023), where unlearning is formulated as reconstructing a decision boundary that could be reasonably learnt by a model trained without the forget data, achieved through learning the nearest false label for each forget sample.
@@
In this work, we approach ZS unlearning from an information theoretic perspective. Golatkar et al. (2020) consider information leakage when observing model weights, whereas we consider the information gained by a model by training on a given sample. Data points offer a classifier different amounts of information gain when included in the training data (Lindley, 1956; Houlsby et al., 2011). If a data point can be inferred from other training data, then it offers little information gain (Jeong and Qiu, 2018). In unlearning, knowing the contribution a sample has made to a model is hard. We postulate that the information gain of a sample can serve as a good indicator for how much influence it has over a model. From this hypothesis, we derive a principled loss to induce forgetting based on how the geometry of the problem space changes with respect to the softmax classifier. If a sample offers little information gain then it may be inferred from other data, therefore the classifier’s output should change minimally over similar samples. This minimal change can be effectively measured via the curvature of the model should be low in the region surrounding such a point. In contrast, a high information gain sample cannot be inferred, and therefore one would expect they lie in high curvature regions of space. We present Just in Time (JiT) unlearning, a novel ZS unlearning algorithm based on minimising the gradients of a classifier with respect to local neighbourhood around each forget sample. We show in low dimensional experiments that removing training samples from low-curvature regions yields minimal changes to the learned decision boundary, whereas removing samples from high curvature regions has a more pronounced effect. We therefore propose that, with reference to Feldman (2020), samples with low information gain may be predicted with more generalised knowledge and thus do not infringe on privacy. In contrast, samples with large information gain have significant impact on the learned classifier, are more likely memorised, and do infringe on privacy. We demonstrate empirically that following these principles, JiT causes the removal of influence from the forget set while preserving generalisation performance across the wider space.
@@ 2 Related Work
Information theory is concerned with the transmission, quantification, and storage of information (Shannon, 1948), and has seen widespread use in machine learning. Most relevant here is its use in determining the information gain of an experiment Lindley (1956). This notion has seen uses such as determining splits in decision trees (Quinlan, 1986), and active learning (Tong and Koller, 2001; Houlsby et al., 2011). Here we use this concept as a proxy to a training sample’s influence on a learned function.
@@
Machine unlearning was first introduced in Cao and Yang (2015), and a probabilistic perspective of unlearning was explored in Ginart et al. (2019); Sekhari et al. (2021); Gupta et al. (2021); Neel et al. (2021); Triantafillou et al. (2023). Here we focus on post-hoc unlearning methods that operate on models that are already trained. Methods exist that alter the initial training scheme (Bourtoule et al., 2021; Mehta et al., 2022; Shah et al., 2023), but these are considered out of scope as they do not satisfy the ZS problem constraints.
@@
Current SOTA methods rely on accessing all or a subset of the original dataset that is not to be forgotten (i.e. the retain set), thus violating the ZS constraints. Bad Teacher unlearning (Chundawat et al., 2023a) and SCRUB (Kurmanji et al., 2023) leverage a student-teacher framework while Amnesiac unlearning (Graves et al., 2021) trains with randomised labels for forget data before fine-tuning on the retained data to repair the model. UNSIR (Tarun et al., 2023) learns an error-maximising noise to induce forgetting of the necessary data, before also employing a finetuning step. Warnecke et al. (2021) minimise the divergence in model output over a sample and its noisy perturbations and then finetune. A key limitation of these methods is that protecting model performance necessitates access to retain-set data for the entire duration of the model’s lifetime. To address this, Golatkar et al. (2020) and Foster et al. (2023) propose methods that do not require fine-tuning or repair steps. Golatkar et al. (2020) derives an unlearning algorithm that minimises information gained about the training data when observing model weights. However, this scales quadratically with dataset size and often performs considerably worse than state-of-the-art (Tarun et al., 2023; Foster et al., 2023). Selective Synaptic Dampening is a scalable retrain-free approach, based on inducing forgetting by selectively dampening parameters that are disproportionately important to the forget-set (Foster et al., 2023). This requires access to the whole dataset at least once, to calculate the importance over the retained data. Chundawat et al. (2023b) introduces two methods to address ZS unlearning. The first method, an extension of Tarun et al. (2023), replaces the repair step with an error-minimising noise. The second utilises a generator network and an attention loss to distil knowledge from an expert teacher, with a band-pass filter preventing the flow of knowledge for specific classes. Both methods are slow, do not scale well to large problem spaces, and can only forget entire classes. Chen et al. (2023) present boundary shrinking and boundary expanding. Shrinking causes unlearning by training over the nearest false label for forget samples, found via a fast gradient sign attack (Goodfellow et al., 2014). While performant, shrinking scales poorly with model and input size. Boundary expanding is faster but less performant, remapping forget samples by training them to fit a new output neuron, before removing the neuron leaving the forget samples in high entropy states.
@@
Membership inference attacks (MIA) are a way of measuring information leakage of a machine learning model (Shokri et al., 2017). An auxiliary model (e.g. a logistic regression) is trained to infer whether a given data point was included in a model’s training data. MIAs are used as typical evaluation metrics in machine unlearning; if a MIA cannot recognise a forgotten sample as an element from the train set, then this presents empirical evidence that the sample has indeed been forgotten.
@@ 3 Preliminaries
We introduce the notation for machine unlearning in a supervised classification task, consistent with the approach outlined in Chen et al. (2023). Consider some input space $\mathcal{X}\subset\mathbb{R}^{d}$ and some output label space $\mathcal{Y}\subset\mathbb{R}^{c}$, where $d$ is the dimensionality of the input and $c$ is the number of classes. We define a training dataset $\mathcal{D}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}^{N}_{i=1}\subseteq\mathcal{X}\times\mathcal{Y}$, where $\boldsymbol{x}_{i}$ is a training input sample with the label $\boldsymbol{y}_{i}$. We denote a subset $\mathcal{D}_{f}\subseteq\mathcal{D}$ as the forget set, and $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{f}$ as the retain set.
@@ 4.2 Entropy Similarity
Figure 3 shows the entropy of the forget-set output distributions for a CNN trained on CIFAR-10. Our unlearning approach increases the entropy over the forget set, reducing the divergence between it and that of a model retrained from scratch on $\mathcal{D}_{r}$. In fact, under a Wilcoxon signed-rank test (Woolson, 2007), we find there is no statistically significant difference between the model unlearned with JiT and the retrained model for $p=0.10$. JiT and retraining both increase the entropy over the forget set, suggesting the resultant models behave in a similar way, possessing less knowledge of the forget samples compared to the baseline model. Alongside matching the entropy over the forget set, JiT preserve model performance, as the unlearned model drops only $2\%$ accuracy (From $99\%$ to $97\%$) on $\mathcal{D}_{r}$. Our algorithm demonstrates promising characteristics that are indicative of an effective unlearning algorithm.
@@ 5.1 Benchmarks
We implement the same benchmarks from Foster et al. (2023), which are similar to that of Chundawat et al. (2023a), Golatkar et al. (2020) and Kurmanji et al. (2023). We run experiments $10$ times, reporting the mean and standard deviation of these performances. Where classes or sub-classes are forgotten, we show performance over the same class/sub-class as in Foster et al. (2023); performance on additional classes can be found in the appendix (10.3). We perform a hyper-parameter search across a single forget class/sub-class, then use these parameters for all classes. This is more realistic, as it cannot be known a priori what future forget sets may be presented to the method. The reported $\mathcal{D}_{r}$ accuracy refers to accuracy over a test set of samples from the classes in $\mathcal{D}_{r}$.
@@
Unlearning scenarios: Typically the three unlearning scenarios are: i) Full-class forgetting, where a full class from the dataset must be unlearned, ii) Sub-class forgetting, where a related subset from a class (e.g. all rockets from class vehicle) is forgotten, and iii) Random forgetting, where a subset is sampled uniformly from the entire training distribution. We evaluate our method in all three scenarios. Datasets: As with previous work, we benchmark JiT on a range of image classification benchmarks. We make use of the CIFAR suite (Krizhevsky and Hinton, 2010), and the Pins Facial Recognition dataset (Burak, 2020), which consists of $17,534$ images of 105 celebrity faces.
@@
Models: We evaluate methods on Vision Transformer (ViT) (Dosovitskiy et al., 2021) and VGG16 (Simonyan and Zisserman, 2014), trained on an NVidia RTX 4090 using Stochastic Gradient Descent with an initial learning rate of $0.1$, and the OneCycle learning rate scheduler (Smith and Topin, 2019).
@@
Evaluation metrics: We evaluate model performance according to four key metrics: i) $\mathcal{D}_{r}$ accuracy, ii) $\mathcal{D}_{f}$ accuracy, iii) MIA score, and iv) method runtime. For all metrics bar runtime, the objective is not to minimise/maximise them, but rather to be as close to the retrained model as possible. This is important, as performing worse than the retrained model implies insufficient performance, but as noted in Chundawat et al. (2023a), Foster et al. (2023) and Kurmanji et al. (2023), significant deviation from the retrained model (e.g. over-forgetting) may leak information about the fact a sample has been forgotten. Finally, there are many forms of membership inference attacks (e.g. LiRA (Carlini et al., 2022)), but to remain consistent with existing unlearning literature we use the same logistic regression MIA as Chundawat et al. (2023a) and Foster et al. (2023).
@@
Comparison methods: We compare JiT against the following methods: i) Baseline (BSLN): that has not been unlearnt, ii) Retrain (RTRN): trained on only the retain data, iii) Finetune (FNTN): , where the model is fine-tuned on $\mathcal{D}_{r}$ for 5 epochs, iv) Selective Synaptic Dampening SSD (Foster et al., 2023), v) GKT (Chundawat et al., 2023b), vi) EMMN (Chundawat et al., 2023b), vii) SCRUB (Kurmanji et al., 2023), viii) Bad Teacher (BT) (Chundawat et al., 2023a) ix) Amnesiac (AMN) (Graves et al., 2021), x) UNSIR (Tarun et al., 2023), xi) Boundary Shrinking (BDSH) (Chen et al., 2023). Since GKT, EMMN, and UNSIR are theoretically limited to forgetting just a full-class, these cannot be evaluated in sub-class or random scenarios. Finally, we note that due to VRAM constraints, we could not benchmark SCRUB on ViT.
@@
JiT hyper-parameters: We conduct a hyper-parameter search for $\eta$ and $\sigma$ using 250 runs of the TPE search from Optuna (Akiba et al., 2019), for each unlearning scenario. For VGG16, we use the following parameters: full-class unlearning uses $\eta=0.0003,\sigma=0.5$, sub-class and random both use $\eta=0.0003,\sigma=0.01$. For ViT, the selected parameters are: full-class $\eta=1.5,\sigma=0.8$, sub-class $\eta=0.5,\sigma=1.5$, and random $\eta=0.01,\sigma=0.5$. ViT and VGG use considerably different learning rates, since only a single epoch is used during the unlearning step. If minimising the runtime is a looser constraint, a smaller learning rate can be used for ViT with extra epochs of training. We stress that when selecting hyper-parameters, we selected values that yielded promising results, without rigorously fitting our results to the retrained model.

### v4

@@ 1 Introduction
Regulations such as the General Data Protection Regulation (GDPR) enshrine an individual’s data autonomy rights, including the right to be forgotten. While deleting an entry from a database is relatively straightforward, removing the influence of that data from a trained model is a challenging open problem. The process of removal, referred to as unlearning, is difficult for several reasons. It is known that neural networks memorise instance-level information (arpit2017closer; zhang2021understanding; feldman2020does), and it is effectively intractable to ascribe parameter changes to a specific training sample post-hoc (kurmanji2023towards). At its core, unlearning is a multi-objective optimization problem with three key desiderata. An effective unlearning algorithm must remove the influence of the selected subset of data, maintain model performance on retained data, and minimise computational cost. These goals are antagonistic since inducing forgetting inevitably disrupts the model’s performance, thus balancing these objectives is key. Naïvely, one may achieve perfect forgetting by retraining a model on the training data sans the forget samples every time there is a forget request. However, this is prohibitively expensive thus violating the third desideratum.
@@
Many of the unlearning methods proposed are effective, however they make strong assumptions about the problem setting that simplifies the task considerably. Primarily, existing methods typically assume access to all, or a subset of, the training data. This data is used in different ways, such as to fine-tune the model post-forgetting (chundawat2023can; graves2021amnesiac), or to conduct parameter importance calculations after the initial training period (foster2023fast). In reality, there are many reasons why this data could be unavailable, such as cost of storage, limited duration access to datasets, or an oversight in considering machine unlearning during model development. As such, chundawat2023zero introduces a novel problem setting for unlearning, termed zero-shot (ZS) unlearning, whereby only the data to be forgotten and the trained model are available. This is extremely challenging, since the remaining data is not available to protect model performance, and thus more delicate methods are required. Insightful treatment of the ZS scenario can be found in chen2023boundary, where unlearning is formulated as reconstructing a decision boundary that could be reasonably learnt by a model trained without the forget data, achieved through learning the nearest false label for each forget sample.
@@
In this work, we approach ZS unlearning from an information theoretic perspective. golatkar2020eternal consider information leakage when observing model weights, whereas we consider the information gained by a model by training on a given sample. Data points offer a classifier different amounts of information gain when included in the training data (lindley1956measure; houlsby2011bayesian). If a data point can be inferred from other training data, then it offers little information gain (jeong2018quantifying). In unlearning, knowing the contribution a sample has made to a model is hard. We postulate that the information gain of a sample can serve as a good indicator for how much influence it has over a model. From this hypothesis, we derive a principled loss to induce forgetting based on how the geometry of the problem space changes with respect to the softmax classifier. If a sample offers little information gain then it may be inferred from other data, therefore the classifier’s output should change minimally over similar samples. This minimal change can be effectively measured via the curvature of the model should be low in the region surrounding such a point. In contrast, a high information gain sample cannot be inferred, and therefore one would expect they lie in high curvature regions of space. We present Just in Time (JiT) unlearning, a novel ZS unlearning algorithm based on minimising the gradients of a classifier with respect to local neighbourhood around each forget sample. We show in low dimensional experiments that removing training samples from low-curvature regions yields minimal changes to the learned decision boundary, whereas removing samples from high curvature regions has a more pronounced effect. We therefore propose that, with reference to feldman2020does, samples with low information gain may be predicted with more generalised knowledge and thus do not infringe on privacy. In contrast, samples with large information gain have significant impact on the learned classifier, are more likely memorised, and do infringe on privacy. We demonstrate empirically that following these principles, JiT causes the removal of influence from the forget set while preserving generalisation performance across the wider space.
@@ 2 Related Work
Information theory is concerned with the transmission, quantification, and storage of information (shannon1948mathematical), and has seen widespread use in machine learning. Most relevant here is its use in determining the information gain of an experiment lindley1956measure. This notion has seen uses such as determining splits in decision trees (quinlan1986induction), and active learning (tong2001support; houlsby2011bayesian). Here we use this concept as a proxy to a training sample’s influence on a learned function.
@@
Machine unlearning was first introduced in cao2015towards, and a probabilistic perspective of unlearning was explored in ginart2019making; sekhari2021remember; gupta2021adaptive; neel2021descent; neurips-2023-machine-unlearning. Here we focus on post-hoc unlearning methods that operate on models that are already trained. Methods exist that alter the initial training scheme (bourtoule2021machine; mehta2022deep; shah2023unlearning), but these are considered out of scope as they do not satisfy the ZS problem constraints.
@@
Current SOTA methods rely on accessing all or a subset of the original dataset that is not to be forgotten (i.e. the retain set), thus violating the ZS constraints. Bad Teacher unlearning (chundawat2023can) and SCRUB (kurmanji2023towards) leverage a student-teacher framework while Amnesiac unlearning (graves2021amnesiac) trains with randomised labels for forget data before fine-tuning on the retained data to repair the model. UNSIR (tarun2023fast) learns an error-maximising noise to induce forgetting of the necessary data, before also employing a finetuning step. warnecke2021machine minimise the divergence in model output over a sample and its noisy perturbations and then finetune. A key limitation of these methods is that protecting model performance necessitates access to retain-set data for the entire duration of the model’s lifetime. To address this, golatkar2020eternal and foster2023fast propose methods that do not require fine-tuning or repair steps. golatkar2020eternal derives an unlearning algorithm that minimises information gained about the training data when observing model weights. However, this scales quadratically with dataset size and often performs considerably worse than state-of-the-art (tarun2023fast; foster2023fast). Selective Synaptic Dampening is a scalable retrain-free approach, based on inducing forgetting by selectively dampening parameters that are disproportionately important to the forget-set (foster2023fast). This requires access to the whole dataset at least once, to calculate the importance over the retained data. chundawat2023zero introduces two methods to address ZS unlearning. The first method, an extension of tarun2023fast, replaces the repair step with an error-minimising noise. The second utilises a generator network and an attention loss to distil knowledge from an expert teacher, with a band-pass filter preventing the flow of knowledge for specific classes. Both methods are slow, do not scale well to large problem spaces, and can only forget entire classes. chen2023boundary present boundary shrinking and boundary expanding. Shrinking causes unlearning by training over the nearest false label for forget samples, found via a fast gradient sign attack (goodfellow2014explaining). While performant, shrinking scales poorly with model and input size. Boundary expanding is faster but less performant, remapping forget samples by training them to fit a new output neuron, before removing the neuron leaving the forget samples in high entropy states.
@@
Membership inference attacks (MIA) are a way of measuring information leakage of a machine learning model (shokri2017membership). An auxiliary model (e.g. a logistic regression) is trained to infer whether a given data point was included in a model’s training data. MIAs are used as typical evaluation metrics in machine unlearning; if a MIA cannot recognise a forgotten sample as an element from the train set, then this presents empirical evidence that the sample has indeed been forgotten.
@@ 3 Preliminaries
We introduce the notation for machine unlearning in a supervised classification task, consistent with the approach outlined in chen2023boundary. Consider some input space $\mathcal{X}\subset\mathbb{R}^{d}$ and some output label space $\mathcal{Y}\subset\mathbb{R}^{c}$, where $d$ is the dimensionality of the input and $c$ is the number of classes. We define a training dataset $\mathcal{D}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}^{N}_{i=1}\subseteq\mathcal{X}\times\mathcal{Y}$, where $\boldsymbol{x}_{i}$ is a training input sample with the label $\boldsymbol{y}_{i}$. We denote a subset $\mathcal{D}_{f}\subseteq\mathcal{D}$ as the forget set, and $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{f}$ as the retain set.
@@ 4.2 Entropy Similarity
Figure 3 shows the entropy of the forget-set output distributions for a CNN trained on CIFAR-10. Our unlearning approach increases the entropy over the forget set, reducing the divergence between it and that of a model retrained from scratch on $\mathcal{D}_{r}$. In fact, under a Wilcoxon signed-rank test (woolson2007wilcoxon), we find there is no statistically significant difference between the model unlearned with JiT and the retrained model for $p=0.10$. JiT and retraining both increase the entropy over the forget set, suggesting the resultant models behave in a similar way, possessing less knowledge of the forget samples compared to the baseline model. Alongside matching the entropy over the forget set, JiT preserve model performance, as the unlearned model drops only $2\%$ accuracy (From $99\%$ to $97\%$) on $\mathcal{D}_{r}$. Our algorithm demonstrates promising characteristics that are indicative of an effective unlearning algorithm.
@@ 5.1 Benchmarks
We implement the same benchmarks from foster2023fast, which are similar to that of chundawat2023can, golatkar2020eternal and kurmanji2023towards. We run experiments $10$ times, reporting the mean and standard deviation of these performances. Where classes or sub-classes are forgotten, we show performance over the same class/sub-class as in foster2023fast; performance on additional classes can be found in the appendix (10.3). We perform a hyper-parameter search across a single forget class/sub-class, then use these parameters for all classes. This is more realistic, as it cannot be known a priori what future forget sets may be presented to the method. The reported $\mathcal{D}_{r}$ accuracy refers to accuracy over a test set of samples from the classes in $\mathcal{D}_{r}$.
@@
Unlearning scenarios: Typically the three unlearning scenarios are: i) Full-class forgetting, where a full class from the dataset must be unlearned, ii) Sub-class forgetting, where a related subset from a class (e.g. all rockets from class vehicle) is forgotten, and iii) Random forgetting, where a subset is sampled uniformly from the entire training distribution. We evaluate our method in all three scenarios. Datasets: As with previous work, we benchmark JiT on a range of image classification benchmarks. We make use of the CIFAR suite (CIFAR_krizhevsky2010convolutional), and the Pins Facial Recognition dataset (pinsds), which consists of $17,534$ images of 105 celebrity faces.
@@
Models: We evaluate methods on Vision Transformer (ViT) (ViT_dosovitskiy2021image) and VGG16 (simonyan2014very), trained on an NVidia RTX 4090 using Stochastic Gradient Descent with an initial learning rate of $0.1$, and the OneCycle learning rate scheduler (smith2019super).
@@
Evaluation metrics: We evaluate model performance according to four key metrics: i) $\mathcal{D}_{r}$ accuracy, ii) $\mathcal{D}_{f}$ accuracy, iii) MIA score, and iv) method runtime. For all metrics bar runtime, the objective is not to minimise/maximise them, but rather to be as close to the retrained model as possible. This is important, as performing worse than the retrained model implies insufficient performance, but as noted in chundawat2023can, foster2023fast and kurmanji2023towards, significant deviation from the retrained model (e.g. over-forgetting) may leak information about the fact a sample has been forgotten. Finally, there are many forms of membership inference attacks (e.g. LiRA (carlini2022membership)), but to remain consistent with existing unlearning literature we use the same logistic regression MIA as chundawat2023can and foster2023fast.
@@
Comparison methods: We compare JiT against the following methods: i) Baseline (BSLN): that has not been unlearnt, ii) Retrain (RTRN): trained on only the retain data, iii) Finetune (FNTN): , where the model is fine-tuned on $\mathcal{D}_{r}$ for 5 epochs, iv) Selective Synaptic Dampening SSD (foster2023fast), v) GKT (chundawat2023zero), vi) EMMN (chundawat2023zero), vii) SCRUB (kurmanji2023towards), viii) Bad Teacher (BT) (chundawat2023can) ix) Amnesiac (AMN) (graves2021amnesiac), x) UNSIR (tarun2023fast), xi) Boundary Shrinking (BDSH) (chen2023boundary). Since GKT, EMMN, and UNSIR are theoretically limited to forgetting just a full-class, these cannot be evaluated in sub-class or random scenarios. Finally, we note that due to VRAM constraints, we could not benchmark SCRUB on ViT.
@@
JiT hyper-parameters: We conduct a hyper-parameter search for $\eta$ and $\sigma$ using 250 runs of the TPE search from Optuna (akiba2019optuna), for each unlearning scenario. For VGG16, we use the following parameters: full-class unlearning uses $\eta=0.0003,\sigma=0.5$, sub-class and random both use $\eta=0.0003,\sigma=0.01$. For ViT, the selected parameters are: full-class $\eta=1.5,\sigma=0.8$, sub-class $\eta=0.5,\sigma=1.5$, and random $\eta=0.01,\sigma=0.5$. ViT and VGG use considerably different learning rates, since only a single epoch is used during the unlearning step. If minimising the runtime is a looser constraint, a smaller learning rate can be used for ViT with extra epochs of training. We stress that when selecting hyper-parameters, we selected values that yielded promising results, without rigorously fitting our results to the retrained model.